At a certain temperature, the equilibrium constant for the following chemical equation is 3.30. SO2(g) + NO2(g) <---> SO3(g) + NO(g), At this temp, calculate the number of moles of NO2(g) that must be added to 2.64 mol of SO2(g) in order to form 1.10 mol of SO3(g) at equilibrium.
Answer #1

K = 3.3

K = SO3*NO/(SO2*NO2)

in equilibrium

[SO2] = 2.64 - x

[NO2] = y - x

[SO3] = 0 + x

[NO] = 0 +x

in equilbirium

[SO3] = 0 + x = 1.1

then

x = 1.1

[SO2] = 2.64 - 1.1 = 1.54

[NO2] = y - 1.1

[SO3] = 0 + 1.1 = 1.1

[NO] = 0 +1.1 = 1.1

solve for y

K = SO3*NO/(SO2*NO2)

3.3 = 1.1*1.1 / (1.54*(y - 1.1 ))

(y-1.1) = 1.1*1.1/(1.54*3.3)

y = 0.2380+1.1 = 1.338 mol of NO2

thats the amount we must add of NO2 in moles